OpenStudy (anonymous):

One solution of the equation 6x2 − 19x + 3 = 0 is 1/6 . What is the other solution?

OpenStudy (jim766):

6x^2 - 19x + 3 do you know how to factor this? or how to start? ( )( )

OpenStudy (anonymous):

no

OpenStudy (anonymous):

can you walk me through it?

OpenStudy (jim766):

6x^2 - 19x + 3 first ask yourself how to mult 2 #s and get 6x^2 (6x )( x ) then ask how you can mult 2 #'s and get 3 (6x 1 )( x 3 )

OpenStudy (jim766):

(6x -1 )( x - 3) = 0

OpenStudy (jim766):

think about when you foil....we are just kind of working it backwards

OpenStudy (jim766):

make any sense at all/

OpenStudy (anonymous):

idk how to foil. but it kinda makes sense

OpenStudy (jim766):

(6x -1)(x -3) = 0 now you can use the zero product poperty 6x-1 = 0 or x -3 = 0 +1 +1 +3 +3 6x = 1 x = 3 x = 1/6